IRR AND NPV

A company is analyzing two mutually exclusive projects, S and L, with the following cash flows:

0 1 2 3 4
Project S -$1,000 $886.38 $250 $15 $10
Project L -$1,000 $0 $240 $400 $843.31

The company's WACC is 9.0%. What is the IRR of the better project? (Hint: The better project may or may not be the one with the higher IRR.) Round your answer to two decimal places.

%
Answer #1

S:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=886.38/1.09+250/1.09^2+15/1.09^3+10/1.09^4

=1042.28

NPV=Present value of inflows-Present value of outflows

=1042.28-1000

=$42.28(Approx).

L:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=240/1.09^2+400/1.09^3+843.31/1.09^4

=1108.30

NPV=Present value of inflows-Present value of outflows

=1108.30-1000

=$108.30(Approx).

Hence L is better having higher NPV.

Let irr be x%
At irr,present value of inflows=present value of outflows.

1000=240/1.0x^2+400/1.0x^3+843.31/1.0x^4

Hence x=irr=12.40%(Approx).

> how did you find irr?

marquez Sun, Nov 21, 2021 11:09 AM